OpenStudy (anonymous):

what number multiplied together gets 36, but added together makes -17

OpenStudy (anonymous):

\(\dfrac{17+\sqrt{145}}{2}\) and \(\dfrac{17-\sqrt{145}}{2}\), I believe.

OpenStudy (anonymous):

My bad...just negate both of those. So: \(-\dfrac{17+\sqrt{145}}{2}\) and \(\dfrac{\sqrt{145}-17}{2}\)